# Analyzing Local Representations of Self-supervised Vision Transformers

### 31 Dec 2023 v1

@@ 5 At Which Granularity Transformers Detect Objects?
Figure 7: Patch retrieval results when the original image is not available. In each subplot, x-axis is the level of blur, y-axis is the number of correctly matched patches in each granularity.
6 How Well ViTs Track Objects?
Object tracking requires identification of the same object instance across frames in a video. In this section we analyze the robustness of patch embeddings over time, as the objects undergo appearance changes.
@@ 6 How Well ViTs Track Objects?
Methodology. We take track-validation set of MOT 2020 images subset of BDD-100K dataset [30] and extract patch representations from each frame. We use ROIAlign [13] to pool features of each object instance based on its ground truth bounding box. For a given time difference $\Delta$ between two frames and for each frame at time step $t$, we retrieve the object with the closest embedding vector from the frame at time step $t+\Delta$. We report the ratio of correctly retrieved instances as the main metric. We also report the percentage of object instances for which the retrieved instance has the same object category. Results are shown in Fig. 8.
DINO-like models are best at tracking. The accuracy of all models degrades for longer time intervals, as appearance changes in average increases over time. The retrieval performance stays relatively constant for intervals larger than one second, i.e. about 30 frames. DINO and DINOv2 show the best performance, with a slight advantage for DINO for $\Delta<10$, which is consistent with the patch retrieval experiments in Sec. 5. Surprisingly, the Supervised ViT is substantially inferior to DINO, and performs slightly better than MAE. We also notice that MAE-200 outperforms MAE across all values of $\Delta$, consistently surpasses Supervised ViT, and achieves comparable accuracy with DINO and DINOv2 for small values of $\Delta$. This indicates that the removal of $200$ high variance features from MAE feature vectors leads to improved performance for tracking.
@@
We repeat the experiments on MOT17 dataset and see a similar behavior. Contrastive approaches strongly outperform ViTs based on masked image modeling. The results are in Sec. A.3.

### 21 Mar 2024 v2

@@ 5.1 Experiments on CityScapes
5.2 Experiments on Fair1M.
For the next set of experiments, we utilized the Fair1M dataset and generated transformed versions of all images. In the first set of experiments, we applied diagonal shifts by $1,2,3,$ and $4$ pixels. In the second set of experiments, we rotated the images by $5,$ $10,$ $15,$ and $20$ degrees counterclockwise. In the third set of experiments, we scaled the images by factors of $0.8$, $0.9$, $1.1$, and $1.2$. For the fourth and fifth sets of experiments, we applied blur and Gaussian noise degradations. We computed patch representations for all these images, creating many images with the same instances of objects.
For the experiments involving image shifting, rotation, and scaling, the first level of evaluation, called ’same-patch’ is challenging. We define the corresponding patch as the one containing the center point of the rotated patch. We remark, that several patches near the corners of the rotated tile may lack a corresponding patch, placing an upper bound on same-patch retrieval accuracy. The results, along with the upper bounds, are shown in Fig. 5. We observe, that for rotation and scaling all models exhibit significantly less robustness even with a 5-degree rotation and a small amount of scaling compared to the highest blur radius or noise level attempted. The order of model performance remains consistent with other image transformations: DINO performs the best, followed by MAE (as well as Scale-MAE) with 200 high-variance features removed, followed by DINOv2 and Supervised ViT, with MAE as a distant outlier. It is worth noting that Scale-MAE outperforms MAE significantly; however, removing its 200 high-variance features also notably improves its performance.
DINO is the most robust one. Among the ViTs we tested, DINO is the most robust one across all transformations. DINOv2 is less robust by all metrics, and performs similar to Supervised ViT. To verify whether this disadvantage of DINOv2 comes from the patch-level loss term or the scale of the model and the dataset, we perform the same analysis with iBOT. iBOT performs slightly better than DINO (Fig. 11 in Appendix), thus patch-level loss cannot be blamed.
Retrieving patches from other images. We notice that most of the retrieved patches come from the same image tile. One potential explanation is that patch embeddings carry image-level information. Repeating the experiment with original image patches excluded, the closest patch can now belong to the same fine-grained category, the same supercategory, an incorrect supercategory, or serve as a background patch. Here DINOv2 takes the lead, Supervised ViT and DINO perform slightly worse, and MAE performs poorly (Fig. 6).
MAE’s performance can be improved. We repeated this experiment with MAE-200 and Scale-MAE-200. These versions of MAE and Scale-MAE perform significantly better than the original ones and outperform both DINOv2 and Supervised ViT, and fall short only of DINO. These results imply that removing the high-variance features of MAE embeddings not only helps in semantic segmentation, but also in the identification of specific object instances in transformed images. On the other hand, this adds to the evidence that the high-variance features do not hold any unique information necessary for identifying the instances, as for small transformations the performance is near ideal  (Fig. 5).
Figure 6: Patch retrieval results when the original image is not available. x-axis is the level of blur, y-axis is the number of correctly matched patches in each granularity.
5.3 Tracking Experiments.
Next, we analyze the robustness of patch embeddings over time, as the objects undergo appearance changes. We report the ratio of correctly retrieved instances and the percentage of object instances for which the retrieved instance has the same object category; see Fig. 7.
@@
DINO-like models are best at tracking. The accuracy of all models degrades for longer time intervals, as appearance changes in average increases over time. The retrieval performance stays relatively constant for intervals larger than one second, i.e. about 30 frames. DINO and DINOv2 show the best performance, with a slight advantage for DINO for $\Delta<10$, which is consistent with the patch retrieval experiments in Sec. 5. Surprisingly, the Supervised ViT is substantially inferior to DINO, and performs slightly better than MAE. We also notice that MAE-200 outperforms MAE across all values of $\Delta$, consistently surpasses Supervised ViT, and achieves comparable accuracy with DINO and DINOv2 for small values of $\Delta$. This indicates that MAE-200 improves the performance of tracking.
@@
We repeat the experiments on MOT17 dataset, observing a consistent pattern: contrastive approaches outperform ViTs based on MIM. Details in Sec. A.3.
6 Discussion
Remedy works for image-level tasks as well: ScaleMAE [27].
To validate our hypothesis, regarding high-variance features affecting distance-based metrics for MAE, we chose a recent extension of MAE called Scale-MAE [27], specifically trained for satellite imagery. Scale-MAE employs k-NN for performance comparison against state-of-the-art methods. We conducted experiments by removing its $m=200$ high-variance features. Scale-MAE is trained in a way that it has a meaningful [CLS] token, which can be used as a representation of an image along with the mean of patch tokens. The authors used [CLS] token representations for image classification using k-NN and showed that it outperforms MAE. We reproduced the results and confirmed that the mean of patch vectors is indeed worse than [CLS]. However, after removing the high-variance features, mean patch representations outperform [CLS] across almost all values of ground sampling distance (GSD) in the UCMerced and RESISC datasets. The results are presented in Fig. 8. Interestingly, the removal of 200 high-variance features from [CLS] does not significantly alter the outcomes.
Figure 8: Comparison between the original Scale-MAE and our proposed variants, where we remove 200 high-variance features, using k-NN on three datasets for various GSDs.
What information do these features hold?
One hypothesis is that these features are necessary for identifying specific instances of objects among the same category, or for distinguishing fine-grained categories of objects. In Sec. 5 we provide negative evidence for this hypothesis: removing high-variance features improves retrieval performance so that the negative effect of losing some instance-specific information, if true, is not detectable.
High variance features do not store global context. Another hypothesis is that the high variance features might contain global information about the image not directly relevant to the semantic class of the current patch. An example of a global information is the number of pixels of each class in the entire image. To test whether patch representations contain such information, we use L2-regularized linear regression to predict those numbers (normalized by the total number of pixels) for the three most common classes (road, building, vegetation) from the representation of a central patch. We report $R^{2}$ for each class and average them. We see that MAE embeddings without top 200 high-variance features have the same predictive power (0.672) as the full MAE embeddings (0.676). It implies that the high variance features do not contain additional information about the global context not present in the other features. Furthermore, the features with the highest variances are less predictive (0.550) than a random subset of 200 features (0.666); see Tab. 6 of Appendix.
Feature normalization has a similar effect for MAE, but not for SimMIM. We created another version of MAE embeddings by applying the pretrained BatchNorm layer from the linear model, before passing them to k-NN classifier. It helped to improve k-NN performance by almost the same amount as removing $m=200$ features. It improved SimMIM’s k-NN performance as well, but linear probing worsened significantly. We conclude that feature normalization is an alternative but not an identical strategy to feature removal.
High variance features of MAE are stable across datasets. We identify top 200 high-variance features of patch representations extracted from Cityscapes, ADE20K and FAIR1M datasets. 196 of these features are shared among Cityscapes and ADE20K, while 192 of them are shared among Cityscapes and FAIR1M.
Distinct layer-wise behavior across ViTs.
We perform a detailed analysis of the k-NN patch classification performance for representations extracted at different layers of the network in Fig. 3. Supervised ViT and DINO perform strikingly similarly. The performance slowly improves from the first layer to the 8th and then saturates. There is a slight drop in performance in the last two layers, which is more noticeable in the case of Supervised ViT. We remark that a similar approach has been observed in [18].
DINOv2’s behavior is quite different. In the first layers its performance is worse than DINO. For the easiest five object categories (road, vegetation, sky, car, building) the performance catches up in the last layer. For harder object categories DINOv2 catches up with DINO and Supervised ViT in the 10th layer, then significantly outperforms them in the 11th and 12th layers. In particular, IoU of bus category jumps from 0.059 on the 9th layer to 0.729 on the 12th layer. The advantage of DINOv2 thus mostly stem from the harder categories. Again, we observe poor k-NN performance in case of the MAE. SimMIM performs better than MAE, but only in the middle layers. The difference is more significant for easier object categories. The quality of the last three layers is similar to MAE.